OpenStudy (anonymous):

Investing. How many years will it take $5000 to grow to $9200 if it is invested at 6% (A) compounded quarterly? (B) compoundedx continuously?

OpenStudy (anonymous):

do you have the formulas for compounded quarterly and continuously? I know I know how to work it, but I don't know the formula off the top of my head.

OpenStudy (tkhunny):

Investing. How many years will it take $5000 to grow to $9200 if it is invested at 6% (A) compounded quarterly? (B) compoundedx continuously? Just write it down! \(5000\left(1 + \dfrac{0.06}{4}\right)^{Yrs\cdot 4} = 9200\) -- Solve for Yrs. Just write it down! \(5000\cdot e^{Yrs\cdot 0.06} = 9200\) -- Solve for Yrs.

OpenStudy (anonymous):

is that the answer? or the formula i need to use?

OpenStudy (anonymous):

what happens to the 9200?

OpenStudy (anonymous):

A) \[5000\times\left( 1.015 \right)^{41}=9206.14\]

OpenStudy (anonymous):

About 41 years for A

OpenStudy (anonymous):

B) \[5000\times e ^{.06\times10.16276} = 9200\]

OpenStudy (anonymous):

exactly 40.9558 years for A. 10.16276 years for B .... hope this helped.

OpenStudy (anonymous):

I actually did that process and i got it wrong?

OpenStudy (anonymous):

Now i have a different amount instead of 9000 i have 17100 for A, instead of 5000 i have 9000 and the percentage stays the same

OpenStudy (tkhunny):

41 is QUARTERS. That's 41/4 = 10.25 Years Any way, Yrs*4*ln(1.015) = ln(9200/5000) and Yrs = 10.23879